DIAL Daily Activity Log Telescope
Date: 7/14/2008   Time Zone: GMT-5 Type: Big Little
Operator: Kuang   Primary Diameter 16'' 4''
Viewing Condition clear  after yesterday's raining F/# 4.5 1
Field Stop 1.5 mrads 4.3mrads
  HV (kV) 1064nm (W) 532nm (W) Yellow(mW) UV(mW) BBO Crystal power after data acquisition Electronics
285 Laser:          65 289.114 49 Ch4 and 5 Ch 1, 2, 3
291 Laser         52 297.025 44 PMT EMI8931QA R7400U-03
found a slight leaking fo r291 last Fri., fix this leaking today and then align the 291 dye laser Voltage 1850V 800V
Are the lasers and tlelescopes aligned before taking data Yes
Ch5 PMT Serial # 49
Ch4 PMT Serial # 28
Primary Activity: Ch1 Ch2 Ch5
Disc Level 8 8 8
Data Description       ADC Range (mV) 100 500
      Filter     Delay (us)   ADC range of Ch1
File Name Time (local) Shots Ch1 Ch2 Ch5 Ch1 Ch2 Ch5  
H0871414.153122 14:14 2400(2min) BP BP BP 0 (constant gate) ~4 0 500
H0871415.563414 15:56 500
Should have used 100mV range for ADC. The saturation below 100m doesn't matter.  
H0871415.590789 100
analyze data, find the channel 5 could be misaligned because ch2 is more consistent with model, then align both lasers with big telescope
H0871416.485345 16:48  
H0871417.491775 17:49  
align 291 again  
H0871505.034466 5:03  
  The data before 18:00 look having problem because channel 1 data/model shape is different with that of channel 2.   


Range-corrected signal

Ozone retrieval (10-min resolution)